# BCQ: Block Clustered Quantization for 4-bit (W4A4) LLM Inference

### v1

@@ Abstract
Post-training quantization (PTQ) is a promising approach to reducing the storage and computational requirements of large language models (LLMs) without additional training cost. Recent PTQ studies have primarily focused on quantizing only weights to sub-$8$-bits while maintaining activations at $8$-bits or higher. Accurate sub-8-bit quantization for both weights and activations without relying on quantization-aware training remains a significant challenge. We propose a novel quantization method called block clustered quantization (BCQ) wherein each operand tensor is decomposed into blocks (a block is a group of contiguous scalars), blocks are clustered based on their statistics, and a dedicated optimal quantization codebook is designed for each cluster. As a specific embodiment of this approach, we propose a PTQ algorithm called Locally-Optimal BCQ (LO-BCQ) that iterates between the steps of block clustering and codebook design to greedily minimize the quantization mean squared error. When weight and activation scalars are encoded to W4A4 format (with $0.5$-bits of overhead for storing scaling factors and codebook selectors), we advance the current state-of-the-art by demonstrating $<1$% loss in inference accuracy across several LLMs and downstream tasks.
Machine Learning, ICML
@@ 1.1 Related work
Recent sub-4-bit quantization proposals such as (Wang et al., 2023; Tseng et al., 2024; Egiazarian et al., 2024) explore extreme weight quantization while maintaining activations at $8$-bit or higher precision. In particular, BitNet (Wang et al., 2023) proposed W1A8 quantization resulting in an aggregate (weights + activations) bitwidth comparable to LO-BCQ. However, BitNet demands training from scratch and despite this large training cost suffers significant loss in accuracy in downstream tasks. QuiP# (Tseng et al., 2024) and AQLM (Egiazarian et al., 2024) propose W2A8 quantization through codebooks. These methods explore vector and additive codebook quantization, respectively, and rely on large codebook sizes ($1$MB of memory footprint) and are not directly applicable to activation quantization. These methods also require weight updates to minimize accuracy loss. In contrast, optimal codebooks ($\leq 0.19$KB of memory footprint) identified by LO-BCQ are applicable to both weight and activation quantization and achieves $<1$% accuracy loss in downstream tasks without any weight updates. Supporting such small codebook sizes makes LO-BCQ more amenable to potential hardware acceleration of decompression. Minimizing quantization MSE using the 1D (Lloyd-Max) and 2D Kmeans clustering has been explored in (Han et al., 2016; Cho et al., 2021, 2023) and (van Baalen et al., 2024), respectively. In contrast, LO-BCQ iteratively optimizes block clustering alongside Lloyd-Max based optimal scalar quantization of block clusters. W4A8 quantization has been proposed in (Frantar et al., 2023; Bai et al., 2021; Yao et al., 2022) involving weight updates to preserve accuracy and in (Lin et al., 2023; van Baalen et al., 2024) without any weight updates (PTQ). Further, (Guo et al., 2023; Wei et al., 2023; Kim et al., 2023a) perform sub-8-bit weight quantization by suppressing outliers.
@@ 1.2 Contributions
  • When each of the weight and activation scalars are quantized to $4$-bits (effective bitwidth including per-block scale-factors etc. is $4.5$ to $4.625$ bits), we achieve $<0.1$ loss in perplexity across GPT3 (1.3B, 8B and 22B) and Llama2 (7B and 70B) models and $<0.2$ loss in Nemotron4 (15B and 340B) models, respectively, on the Wikitext-103 dataset. Further, we achieve $<1$% loss in average accuracy across downstream tasks such as MMLU and LM evaluation harness.
To the best of our knowledge, we are the first to achieve $<1$% loss in downstream task accuracy when both LLM activations and weights are quantized to $4$-bits during PTQ (no fine-tuning or weight updates).
@@ 2.2 Locally optimal block clustered quantization
Our goal is to construct a family of codebooks $\mathcal{C}$ resulting in minimal quantization MSE during block clustered quantization. Figure 3 presents an algorithm called Locally Optimal BCQ (LO-BCQ) to achieve this goal. LO-BCQ consists of two main steps: (i) updating block clusters with fixed per-cluster codebooks, and (ii) updating per-cluster codebooks with fixed block clusters. This algorithm begins at iteration 00 (initial condition) with a set of $N_{c}$ initial codebooks $\{C_{1}^{(0)},\ldots,C_{N_{c}}^{(0)}\}$ and unquantized operand blocks as inputs. During step 1 of iteration $n$, with the per-cluster codebooks from the previous iteration $\{C_{1}^{(n-1)},\ldots C_{N_{c}}^{(n-1)}\}$, we perform block clustering by mapping each block to the codebook that achieves minimum quantization MSE. That is, we use the following mapping function:
@@ 2.4 Block formats for LO-BCQ
Figure 5 illustrates the LO-BCQ block format where each operand block of length $L_{b}$ is associated with a $log2(N_{c})$-bit index (result of the mapping function $f$ in equation 4) that selects the best codebook for that block. Each codebook is composed of $2^{B}$ entries and each scalar in the operand block is a $B$-bit index that represents the closest value in the selected codebook. Each entry in the codebook is a $B_{c}$ [^3] We set $B_{c}=6$ in practice-bit integer. Finally, each block array $\bm{A}$ is associated with a scale-factor $s_{A}$. This scale-factor and its quantization $\hat{s}_{A}$ to $B_{s}$ [^4] We set $B_{s}=8$ and the scale factor format is E4M3-bits are computed as:
@@
where the term ${N_{c}*2^{B}*B_{c}}/{L_{X}}$ is usually negligible since the memory footprint of codebooks (numerator) is negligible compared to the size of the operand tensor (denominator). Indeed, we emphasize that LO-BCQ shares a set of $N_{c}<=16$ codebooks of size $<=0.19KB$ among the scalars of the entire tensor, resulting in negligible memory overhead for storing the codebooks.
Figure 6: (left) LO-BCQ codebooks compared to $4$-bit floating point formats (center) Layerwise normalized MSE (NMSE) compared to $4$-bit block formats. We compute NMSE for the weights of first $20$ GEMM layers (QKV, projection and fully-connected) of Llama2-7B model. Note that we use the NMSE for better visualization across varying layer data. (right) Quantization NMSE acheived by universally calibrated codebooks compared to that calibrated layerwise in Llama2-7B inputs of first $30$ GEMM (QKV, projection and fully-connected) layers.
@@ 3 Applying LO-BCQ for LLM Inference
Figure 6 (left) compares the codebooks identified by the LO-BCQ algorithm in a GEMM layer of a GPT3-126M model to $4$-bit floating point formats such as E1M2, E2M1 and E3M0. The LO-BCQ codebooks outperform other block formats as shown in Figure 6 (center) by capturing the arbitrary and non-uniform patterns in the value distributions of LLM operands and allowing each block to map to the codebook that best represents it. The mapping of operand blocks to the best of available codebooks can be conceptually compared to prior works that have explored mixed-format quantization such as (Tambe et al., 2020; Zadeh et al., 2022).
LO-BCQ provides the quantization operation the flexibility to assign data to any of the sign posts (codewords) in Figure 6(left). The union of these sign posts covers the real line with a resolution that is clearly superior to that of a 4-bit quantizer. Therefore, we hypothesized that these codebooks need not be calibrated on a per-tensor (layerwise) basis, but rather, it is likely that they would be universally appropriate to quantize any tensor (weights and activations), at any layer, for any model. To verify this hypothesis, we calibrated a set of codebooks on data sampled from GPT3 models on Wikitext-103 dataset and froze it. We find that these codebooks achieve comparable quantization MSE compared to those calibrated individually on each operand as shown in Figure 6 (right) which verifies our hypothesis. In our subsequent results, we always employ universally calibrated codebooks.
Finally, the small size of LO-BCQ codebooks enables efficient activation quantization on the fly. Indeed, LO-BCQ involves computing the following values – per-block array scale-factor $s_{A}$, per-block codebook selector $s_{b}$ which is the result of the mapping function $f$ (Eq. 4), and the index to closest representation $\hat{b}$ in the selected codebook (Eq. 2). Note that the computation of $s_{A}$, similar to other block quantization methods (Rouhani et al., 2023b; Dai et al., 2021), simply corresponds to a max-reduction (followed by quantization) over the block array whose size is small ($<=128$). Further, with LO-BCQ, the codebooks are constant (frozen) with small size ($<=0.19KB$). This is an important distinction with other works on codebook quantization (Tseng et al., 2024; Egiazarian et al., 2024). As such, $s_{b}$ and $\hat{b}$ can be concurrently computed across operand blocks.
@@ 4 Experimental Evaluation of LO-BCQ
Method Bitwidth Wikitext-103 PPL ($\Delta$)
(W4A4) GPT3 Llama2 Nemotron4
1.3B 8B 22B 7B 70B 15B 340B
BF16 (Pretrained) 16 9.98 7.38 6.54 5.06 3.14 5.87 3.48
MX4 (g16) 4.5 11.33 (1.35) 8.15 (0.77) 7.69 (1.15) 5.73 (0.67) 3.58 (0.44) 8.88 (3.01) 4.01 (0.53)
VSQ (g16) 4.5 10.83 (0.85) 8.17 (0.79) 7.12 (0.58) 835 (829) 4.96 (1.82) 7.58 (1.71) 4.19 (0.71)
MXFP4 (g32) 4.25 11.04 (1.06) 9.12 (1.74) 10.18 (3.64) 5.76 (0.70) 3.69 (0.55) 8.24 (2.37) 4.10 (0.62)
LO-BCQ (g64, 2 codebooks) 4.25 10.40 (0.42) 7.61 (0.23) 6.74 (0.20) 5.31 (0.25) 3.35 (0.21) 6.30 (0.43) 3.67 (0.19)
LO-BCQ (g64, 8 codebooks) 4.5 10.17 (0.19) 7.48 (0.10) 6.62 (0.08) 5.19 (0.13) 3.23 (0.09) 6.13 (0.26) 3.60 (0.12)
LO-BCQ (g32, 16 codebooks) 4.75 10.12 (0.14) 7.45 (0.07) 6.59 (0.05) 5.15 (0.09) 3.20 (0.06) 6.03 (0.16) 3.56 (0.08)
@@ 4.1 Experimental Setup
We perform accuracy studies on GPT3 (Shoeybi et al., 2020) (1.3B, 8B and 22B), Llama2 (Touvron et al., 2023) (7B and 70B) and Nemotron4 (15B and 340B) (Parmar et al., 2024) models. We evaluate PTQ inference accuracy on several downstream tasks including Wikitext-103 (Merity et al., 2016), MMLU (Hendrycks et al., 2021) and Eleuther AI’s LM evaluation harness (Gao et al., 2024). In LM evaluation harness, we infer on Race (RA), Boolq (BQ), Hellaswag (HS), Piqa (PQ) and Winogrande (WG) tasks and in the MMLU dataset we evaluate all tasks. In all these models, we quantize GEMM layers including Query, Key and Value computations, Projection layer after self attention and the fully-connected layers.
@@ 4.2 Accuracy studies on downstream tasks
We present our comprehensive accuracy evaluations across the Nemotron4, Llama2 and GPT3 models, on the Wikitext-103, LM evaluation harness and MMLU datasets. For convenience, we present select LO-BCQ configurations with $L_{b}=8$ in this section. See 10 for accuracy studies on other configurations.
@@ 4.2.1 Perplexity on Wikitext-103
.   Method Bitwidth $\Delta$ Wiki PPL (g128) (W4A4) Llama2-7B Llama2-70B SmoothQuant 4.13 77.65 - OmniQuant 4.13 9.14 - QuaRot 4.13 0.46 0.29 Atom 4.13 0.56 0.36 LO-BCQ (2 codebooks) 4.19 0.14 0.09 LO-BCQ (4 codebooks) 4.31 0.12 0.07 LO-BCQ (8 codebooks) 4.44 0.09 0.06 LO-BCQ (16 codebooks) 4.56 0.08 0.05
@@
Method Bitwidth Llama2-7B Llama2-70B
(W4A4) RA BQ WG PQ HS Avg ($\Delta$) RA BQ WG PQ HS Avg ($\Delta$)
BF16 16 44.4 79.29 69.38 78.07 57.10 65.65 48.8 85.23 79.95 81.56 65.27 72.16
QuaRot (g128) 4.13 - - 63.77 76.77 - - - - 76.24 82.43 - -
MX4 (g16) 4.5 41.43 73.98 66.22 77.04 55.19 62.77 (2.88) 48.04 82.41 76.40 80.58 63.24 70.13 (2.03)
VSQ (g16) 4.5 31.39 65.75 55.49 67.30 43.51 52.69 (12.96) 47.85 82.29 77.27 79.82 61.40 69.73 (2.43)
MXFP4 (g32) 4.25 41.34 74.00 67.48 77.53 54.22 62.91 (2.74) 47.75 83.06 76.32 80.58 63.24 70.19 (1.97)
LO-BCQ (g64, 2 codebooks) 4.25 42.49 77.58 68.90 77.09 55.93 64.40 (1.25) 49.0 82.82 78.77 81.45 64.21 71.25 (0.91)
LO-BCQ (g64, 8 codebooks) 4.5 42.58 77.43 69.77 77.09 56.51 64.68 (0.97) 49.28 84.03 78.37 81.45 64.76 71.58 (0.58)
LO-BCQ (g32, 16 codebooks) 4.75 43.73 77.86 68.90 77.86 56.52 64.97 (0.68) 49.28 84.93 80.66 81.34 65.18 72.28 (+0.12)
GPT3-8B GPT3-22B
BF16 16 41.34 68.32 67.88 78.78 54.16 62.10 40.67 76.54 70.64 79.16 57.11 64.82
MX4 (g16) 4.5 38.28 66.27 65.11 75.63 50.77 59.21 (2.89) 39.04 72.26 67.96 77.86 54.77 62.38 (2.44)
VSQ (g16) 4.5 40.86 63.91 66.93 76.28 51.38 59.87 (2.23) 40.57 65.81 69.61 77.20 54.82 61.60 (3.22)
MXFP4 (g32) 4.25 39.71 65.35 67.01 76.12 50.22 59.68 (2.42) 39.14 69.61 64.17 75.68 47.60 59.24 (5.58)
LO-BCQ (g64, 2 codebooks) 4.25 40.48 69.20 66.85 77.31 53.06 61.38 (0.72) 40.48 75.41 69.14 78.24 56.06 63.87 (0.95)
LO-BCQ (g64, 8 codebooks) 4.5 39.43 69.45 67.72 77.75 53.71 61.61 (0.49) 39.43 77.09 70.17 78.62 56.60 64.38 (0.44)
LO-BCQ (g32, 16 codebooks) 4.75 39.62 69.30 67.00 77.37 53.51 61.36 (0.74) 39.62 75.35 69.30 78.89 56.64 63.96 (0.86)

### v2

# LO-BCQ: Locally Optimal Block Clustered Quantization for 4-bit (W4A4) LLM Inference

@@ Abstract
Post-training quantization (PTQ) is a promising approach to reducing the storage and computational requirements of large language models (LLMs) without additional training cost. Recent PTQ studies have primarily focused on quantizing only weights to sub-$8$-bits while maintaining activations at $8$-bits or higher. Accurate sub-8-bit quantization for both weights and activations without relying on quantization-aware training remains a significant challenge. We propose a novel quantization method called block clustered quantization (BCQ) wherein each operand tensor is decomposed into blocks (a block is a group of contiguous scalars), blocks are clustered based on their statistics, and a dedicated optimal quantization codebook is designed for each cluster. As a specific embodiment of this approach, we propose a PTQ algorithm called Locally-Optimal BCQ (LO-BCQ) that iterates between the steps of block clustering and codebook design to greedily minimize the quantization mean squared error. When weight and activation scalars are encoded to W4A4 format (with $0.5$-bits of overhead for storing scaling factors and codebook selectors), we advance the current state-of-the-art by demonstrating <1<1% loss in inference accuracy across several LLMs and downstream tasks.
@@ 1.1 Related work
[…]
Recent sub-4-bit quantization proposals such as (Wang et al., 2023; Tseng et al., 2024; Egiazarian et al., 2024) explore extreme weight quantization while maintaining activations at $8$-bit or higher precision. In particular, BitNet (Wang et al., 2023) proposed W1A8 quantization resulting in an aggregate (weights + activations) bitwidth comparable to LO-BCQ. However, BitNet demands training from scratch and despite this large training cost suffers significant loss in accuracy in downstream tasks. QuiP# (Tseng et al., 2024) and AQLM (Egiazarian et al., 2024) propose W2A8 quantization through codebooks. These methods explore vector and additive codebook quantization, respectively, and rely on large codebook sizes (11MB of memory footprint) and are not directly applicable to activation quantization. These methods also require weight updates to minimize accuracy loss. In contrast, optimal codebooks ($\leq 0.19$KB of memory footprint) identified by LO-BCQ are applicable to both weight and activation quantization and achieves <1<1% accuracy loss in downstream tasks without any weight updates. Supporting such small codebook sizes makes LO-BCQ more amenable to potential hardware acceleration of decompression. Minimizing quantization MSE using the 1D (Lloyd-Max) and 2D Kmeans clustering has been explored in (Han et al., 2016; Cho et al., 2021; 2023) and (van Baalen et al., 2024), respectively. In contrast, LO-BCQ iteratively optimizes block clustering alongside Lloyd-Max based optimal scalar quantization of block clusters. W4A8 quantization has been proposed in (Frantar et al., 2023; Bai et al., 2021; Yao et al., 2022) involving weight updates to preserve accuracy and in (Lin et al., 2023; van Baalen et al., 2024) without any weight updates (PTQ). Further, (Guo et al., 2023; Wei et al., 2023; Kim et al., 2023a) perform sub-8-bit weight quantization by suppressing outliers.
@@ 1.2 Contributions
  • When each of the weight and activation scalars are quantized to $4$-bits (effective bitwidth including per-block scale-factors etc. is $4.5$ to $4.625$ bits), we achieve <0.1<0.1 loss in perplexity across GPT3 (1.3B, 8B and 22B) and Llama2 (7B and 70B) models and <0.2<0.2 loss in Nemotron4 (15B and 340B) models, respectively, on the Wikitext-103 dataset. Further, we achieve <1<1% loss in average accuracy across downstream tasks such as MMLU and LM evaluation harness.
To the best of our knowledge, we are the first to achieve <1<1% loss in downstream task accuracy when both LLM activations and weights are quantized to $4$-bits during PTQ (no fine-tuning or weight updates).
@@ 2.1 Mathematical Definition
In practice, groups of blocks called a block-array share $8$-bit scale factors each and the codebook entries are $6$-bit integers. We discuss this further in section 2.4
@@ 2.2 Locally optimal block clustered quantization
Our goal is to construct a family of codebooks $\mathcal{C}$ resulting in minimal quantization MSE during block clustered quantization. Figure 3 presents an algorithm called Locally Optimal BCQ (LO-BCQ) to achieve this goal. LO-BCQ consists of two main steps: (i) updating block clusters with fixed per-cluster codebooks, and (ii) updating per-cluster codebooks with fixed block clusters. This algorithm begins at iteration 0 (initial condition) with a set of $N_{c}$ initial codebooks $\{C_{1}^{(0)},\ldots,C_{N_{c}}^{(0)}\}$ and unquantized operand blocks as inputs. During step 1 of iteration $n$, with the per-cluster codebooks from the previous iteration $\{C_{1}^{(n-1)},\ldots C_{N_{c}}^{(n-1)}\}$, we perform block clustering by mapping each block to the codebook that achieves minimum quantization MSE. That is, we use the following mapping function:
@@ 2.4 Block formats for LO-BCQ
Figure 5 illustrates the LO-BCQ block format where each operand block of length $L_{b}$ is associated with a $log2(N_{c})$-bit index (result of the mapping function $f$ in equation 4) that selects the best codebook for that block. Each codebook is composed of $2^{B}$ entries and each scalar in the operand block is a $B$-bit index that represents the closest value in the selected codebook. Each entry in the codebook is a $B_{c}$-bit integer. Finally, each block array $\bm{A}$ is associated with a scale-factor $s_{A}$. This scale-factor and its quantization $\hat{s}_{A}$ to $B_{s}$-bits are computed as:
@@
where the term ${N_{c}*2^{B}*B_{c}}/{L_{X}}$ is usually negligible since the memory footprint of codebooks (numerator) is negligible compared to the size of the operand tensor (denominator). Indeed, we emphasize that LO-BCQ shares a set of $N_{c}<=16$ codebooks of size <=0.19KB<=0.19KB among the scalars of the entire tensor, resulting in negligible memory overhead for storing the codebooks.
In this paper, we assume $B_{s}=8$ and the data format $F$ is floating point E4M3. Further, each codebook entry is a $6$-bit integer (i.e, $B_{c}=6$) and we vary $N_{c}$ between $2$ and $16$, $L_{b}$ between $2$ and $8$, and $L_{A}$ between $16$ and $128$ to obtain various LO-BCQ configurations. We list the configurations and their corresponding bitwidths in Table 1. We perform a detailed ablation of these configurations and present our insights in section 4.3.
Table 1: Various LO-BCQ configurations and their bitwidths. Each block array of size $L_{A}$ share an $8$-bit scale factor. Each operand block of size $L_{b}$ is associated with a $log2(N_{c})$-bit index that maps it to one out of $N_{c}$ codebooks that best represents it.
$L_{b}=8$ $L_{b}=4$ $L_{b}=2$
$L_{A}$ $N_{c}$ 2 4 8 16 2 4 2
128 4.1875 4.3125 4.4375 4.5625 4.3125 4.5625 4.5625
64 4.25 4.375 4.5 4.625 4.375 4.625 4.625
32 4.375 4.5 4.625 4.75 4.5 4.75 4.75
16 4.625 4.75 4.875 5 4.75 5 5
Figure 6: LO-BCQ codebooks compared to $4$-bit floating point formats and layerwise normalized MSE (NMSE). We compute NMSE for the weights of first 2020 GEMM layers (QKV, projection and fully-connected) of Llama2-7B model. Note that we use the NMSE for better visualization across varying layer data.
@@ 3 Applying LO-BCQ for LLM Inference
Figure 6 (left) compares the codebooks identified by the LO-BCQ algorithm in a GEMM layer of a GPT3-126M model to $4$-bit floating point formats such as E1M2, E2M1 and E3M0. The LO-BCQ codebooks outperform other block formats as shown in Figure 6 (right) by capturing the arbitrary and non-uniform patterns in the value distributions of LLM operands and allowing each block to map to the codebook that best represents it. The mapping of operand blocks to the best of available codebooks can be conceptually compared to prior works that have explored mixed-format quantization such as (Tambe et al., 2020; Zadeh et al., 2022).
LO-BCQ provides the quantization operation the flexibility to assign data to any of the sign posts (codewords) in Figure 6(left). The union of these sign posts covers the real line with a resolution that is clearly superior to that of a 4-bit quantizer. Therefore, we hypothesized that these codebooks need not be calibrated on a per-tensor (layerwise) basis, but rather, it is likely that they would be universally appropriate to quantize any tensor (weights and activations), at any layer, for any model. To verify this hypothesis, we calibrated a set of codebooks on data sampled from GPT3 models on Wikitext-103 dataset and froze it. We find that these codebooks achieve comparable quantization MSE compared to those calibrated individually on each operand as shown in Figure 7 which verifies our hypothesis. In our subsequent results, we always employ universally calibrated codebooks.
Figure 7: Quantization NMSE acheived by universally calibrated codebooks compared to that calibrated layerwise in Llama2-7B inputs of first $30$ GEMM (QKV, projection and fully-connected) layers.
Finally, the small size of LO-BCQ codebooks enables efficient activation quantization on the fly. Indeed, LO-BCQ involves computing the following values – per-block array scale-factor $s_{A}$, per-block codebook selector $s_{b}$ which is the result of the mapping function $f$ (Eq. 4), and the index to closest representation $\hat{b}$ in the selected codebook (Eq. 2). Note that the computation of $s_{A}$, similar to other block quantization methods (Rouhani et al., 2023b; Dai et al., 2021), simply corresponds to a max-reduction (followed by quantization) over the block array whose size is small (<=128<=128). Further, with LO-BCQ, the codebooks are constant (frozen) with small size (<=0.19KB<=0.19KB). This is an important distinction with other works on codebook quantization (Tseng et al., 2024; Egiazarian et al., 2024). As such, $s_{b}$ and $\hat{b}$ can be concurrently computed across operand blocks.
@@ 4.1 Experimental Setup
We perform accuracy studies on GPT3 (Shoeybi et al., 2020) (1.3B, 8B and 22B), Llama2 (Touvron and others, 2023) (7B and 70B) and Nemotron4 (15B and 340B) (Parmar et al., 2024) models. We evaluate PTQ inference accuracy on several downstream tasks including Wikitext-103 (Merity et al., 2016), MMLU (Hendrycks et al., 2021) and Eleuther AI’s LM evaluation harness (Gao et al., 2024). In LM evaluation harness, we infer on Race (RA), Boolq (BQ), Hellaswag (HS), Piqa (PQ) and Winogrande (WG) tasks and in the MMLU dataset we evaluate all tasks. In all these models, we quantize GEMM layers including Query, Key and Value computations, Projection layer after self attention and the fully-connected layers.
@@ 4.2 Accuracy studies on downstream tasks
We present our comprehensive accuracy evaluations across the Nemotron4, Llama2 and GPT3 models, on the Wikitext-103, LM evaluation harness and MMLU datasets. For convenience, we present select LO-BCQ configurations with $L_{b}=8$ in this section. See A.3 for accuracy studies on other configurations.
@@ 4.2.1 Perplexity on Wikitext-103
Method Bitwidth Wikitext-103 PPL ($\Delta$)
(W4A4) GPT3 Llama2 Nemotron4
8B 22B 7B 70B 15B 340B
BF16 (Pretrained) 16 7.38 6.54 5.06 3.14 5.87 3.48
MX4 (g16) 4.5 8.15 (0.77) 7.69 (1.15) 5.73 (0.67) 3.58 (0.44) 8.88 (3.01) 4.01 (0.53)
VSQ (g16) 4.5 8.17 (0.79) 7.12 (0.58) 835 (829) 4.96 (1.82) 7.58 (1.71) 4.19 (0.71)
MXFP4 (g32) 4.25 9.12 (1.74) 10.18 (3.64) 5.76 (0.70) 3.69 (0.55) 8.24 (2.37) 4.10 (0.62)
LO-BCQ (g64, $N_{c}=2$) 4.25 7.61 (0.23) 6.74 (0.20) 5.31 (0.25) 3.35 (0.21) 6.30 (0.43) 3.67 (0.19)
LO-BCQ (g64, $N_{c}=8$) 4.5 7.48 (0.10) 6.62 (0.08) 5.19 (0.13) 3.23 (0.09) 6.13 (0.26) 3.60 (0.12)
LO-BCQ (g32, $N_{c}=16$) 4.75 7.45 (0.07) 6.59 (0.05) 5.15 (0.09) 3.20 (0.06) 6.03 (0.16) 3.56 (0.08)
@@
.   Method Bitwidth $\Delta$ Wiki PPL (g128) (W4A4) Llama2-7B Llama2-70B SmoothQuant 4.13 77.65 - OmniQuant 4.13 9.14 - QuaRot 4.13 0.46 0.29 Atom 4.13 0.56 0.36 LO-BCQ ($N_{c}=2$) 4.19 0.14 0.09 LO-BCQ ($N_{c}=4$) 4.31 0.12 0.07 LO-BCQ ($N_{c}=8$) 4.44 0.09 0.06 LO-BCQ ($N_{c}=16$) 4.56 0.08 0.05
@@
Table 5: Comparing perplexity loss of sub-4-bit weight-only LO-BCQ to other weight-only quantization methods.
Method Bitwidth #codebooks Wiki2 PPL
Llama2-7B Llama2-13B Llama2-70B
BF16 (Pretrained) 16 - 5.47 4.88 3.32
W3A16 QuIP# (LDLQ, no FT) 3 $2^{16}+2^{8}$ 5.91 5.23 3.61
AQLM (FT) 3 $2*2^{12}$ 5.46 4.82 3.36
3.375 4 5.79 5.12 3.53
LO-BCQ (LDLQ, no FT) 3.5 8 5.72 5.09 3.49
W2A16 QuIP# (LDLQ, no FT) 2 $2^{16}$ 8.05 6.59 4.44
AQLM (FT) 2 $2^{16}$ 6.59 5.60 3.94
2.375 4 8.02 7.12 4.52
LO-BCQ (LDLQ, no FT) 2.5 8 6.87 6.16 4.20
@@
Method Bitwidth Llama2-7B
(W4A4) RA BQ WG PQ HS Avg ($\Delta$ %)
BF16 16 44.4 79.29 69.38 78.07 57.10 65.65
QuaRot (g128) 4.13 - - 63.77 76.77 - -
MX4 (g16) 4.5 41.43 73.98 66.22 77.04 55.19 62.77 (2.88)
VSQ (g16) 4.5 31.39 65.75 55.49 67.30 43.51 52.69 (12.96)
MXFP4 (g32) 4.25 41.34 74.00 67.48 77.53 54.22 62.91 (2.74)
LO-BCQ (g64, $N_{c}=2$) 4.25 42.49 77.58 68.90 77.09 55.93 64.40 (1.25)
LO-BCQ (g64, $N_{c}=8$) 4.5 42.58 77.43 69.77 77.09 56.51 64.68 (0.97)
LO-BCQ (g32, $N_{c}=16$) 4.75 43.73 77.86 68.90 77.86 56.52 64.97 (0.68)
Llama2-70B
BF16 16 48.8 85.23 79.95 81.56 65.27 72.16
QuaRot (g128) 4.13 - - 76.24 82.43 - -
MX4 (g16) 4.5 48.04 82.41 76.40 80.58 63.24 70.13 (2.03)
VSQ (g16) 4.5 47.85 82.29 77.27 79.82 61.40 69.73 (2.43)
MXFP4 (g32) 4.25 47.75 83.06 76.32 80.58 63.24 70.19 (1.97)
LO-BCQ (g64, $N_{c}=2$) 4.25 49.0 82.82 78.77 81.45 64.21 71.25 (0.91)
LO-BCQ (g64, $N_{c}=8$) 4.5 49.28 84.03 78.37 81.45 64.76 71.58 (0.58)
LO-BCQ (g32, $N_{c}=16$) 4.75 49.28 84.93 80.66 81.34 65.18 72.28 (+0.12)
GPT3-8B
BF16 16 41.34 68.32 67.88 78.78 54.16 62.10
MX4 (g16) 4.5 38.28 66.27 65.11 75.63 50.77 59.21 (2.89)
VSQ (g16) 4.5 40.86 63.91 66.93 76.28 51.38 59.87 (2.23)
MXFP4 (g32) 4.25 39.71 65.35 67.01 76.12 50.22 59.68 (2.42)
LO-BCQ (g64, $N_{c}=2$) 4.25 40.48 69.20 66.85 77.31 53.06 61.38 (0.72)
LO-BCQ (g64, $N_{c}=8$) 4.5 39.43 69.45 67.72 77.75 53.71 61.61 (0.49)
LO-BCQ (g32, $N_{c}=16$) 4.75 39.62 69.30 67.00 77.37 53.51 61.36 (0.74)
GPT3-22B
BF16 16 40.67 76.54 70.64 79.16 57.11 64.82
MX4 (g16) 4.5 39.04 72.26 67.96 77.86 54.77 62.38 (2.44)
VSQ (g16) 4.5 40.57 65.81 69.61 77.20 54.82 61.60 (3.22)
MXFP4 (g32) 4.25 39.14 69.61 64.17 75.68 47.60 59.24 (5.58)
LO-BCQ (g64, $N_{c}=2$) 4.25 40.48 75.41 69.14 78.24 56.06 63.87 (0.95)
LO-BCQ (g64, $N_{c}=8$) 4.5 39.43 77.09 70.17 78.62 56.60 64.38 (0.44)
LO-BCQ (g32, $N_{c}=16$) 4.75 39.62 75.35 69.30 78.89 56.64 63.96 (0.86)